Faye is 31 years older than her daughter Ilana, but 2 years younger than her husband Leo. If the sum of their three ages is 100, how old is Ilana?
Is someone able to explain how I go about working this figure out please? Many thanks.

Answer by kyawkhant(2)

Solution:
Let Fayae age be "x" year
Then, Ilana age = x-31
Leo age = x+2
The sumof their three ages is 100
x+x-31+x+2=100
3x-29=100
3x=129
x=43
Ilana age= x-31
= 43-31
=12 years old.